# EXCGEC: A Benchmark of Edit-wise Explainable Chinese Grammatical Error Correction

### v1

@@ 5.2 Inference
For post-explaining EXGEC models, we design a specific Correct-Then-Explain decoding algorithm called COTE, which is presented in Algorithm 1. First, we employ the greedy beam search decoding strategy for the correction part, which is beneficial to relieve the over-correction problem that is common on LLMs Cao et al. (2023); Loem et al. (2023); Li et al. (2023a). Then, we apply CLEME to extract edits. Notably, we merge adjacent edits with distance less than 2 characters to avoid fragmented edits. Finally, we leverage the Top-p decoding strategy for generating explanations, encouraging diversified natural language explanations. It is worth noting that COTE is not accessible to pre-explaining models since the edit extraction tool necessitates both a source text and a target text.
6 Experiments
@@ 6.1 Experimental Settings
We benchmark three series of LLMs, including Llama-3 Touvron et al. (2023), Qwen-1.5 Bai et al. (2023), and DeepSeek Bi et al. (2024). For each series of LLMs, we experiment with their base and chat (or instruct) versions to investigate whether further alignment training benefits the task. All results are based on EXCGEC-test. Training details are reported in Appendix B.1.
@@
We report experiment results using the metrics introduced in Section 4.4, calculated using open-source toolkits including NLTK Bird and Loper (2004), rouge Lin (2004), and scikit-learn Pedregosa et al. (2011). Particularly, we observe many hypothesis edits are not covered in references, making it impossible to evaluate the subsequent explanations for these edits. To address this, we introduce two extra indicators, namely Hit and Miss rates. A hypothesis edit overlapping with a reference edit is designated as a hit edit, while a reference edit without any match with hypothesis edits is deemed a miss edit. The hit rate is defined as the ratio of hit edits to all hypothesis edits, and the miss rate as the ratio of miss edits to all reference edits. Only hit edits are used to determine the evaluation outcomes for explanations.
@@
6.2 Results of Multi-task Models
The preliminary results from both the post- and pre-explaining models are presented in Table 3, from which we can make some conclusions.
Post-explaining models consistently outperform pre-explaining models.
In relation to the correction aspect, all post-explaining models obtain higher F0.5 scores than pre-explaining models, regardless of the applied backbones. A similar pattern is observed in the explanation part, where all the pre-explaining models invariably underperform their post-explaining counterparts. This suggests that a complexity for LLMs in initially explaining grammatical errors. And once pre-explaining models generate flawed explanations, the ensuing distraction impedes their ability to accurately correct the source text.
Chat models outperform base models.
For post-explaining models, we observe all chat or instruct models gain slightly higher F0.5 correction scores, and they also marginally outperform their base version counterparts in the explanation task. It indicates that additional alignment training Wang et al. (2023) can benefit the EXGEC task.
@@ 6.2 Results of Multi-task Models
[…]
Beam search 28.31 / 21.21 / 26.54 36.74 / 17.26 / 29.98 99.22 19.05 83.93 44.48 0.71 22.71 44.28 55.55 / 32.26 / 42.34
Top-p 19.45 / 27.05 / 20.61 24.83 / 19.14 / 23.44 99.93 0.40 81.53 39.56 0.74 17.88 41.40 51.73 / 25.81 / 36.51
Table 6: Comparison of the post-explaining model with different token-wise decoding strategies. Note that the explanation performance is conditioned on ground truth target texts in order to exclude unrelated interference.
6.3 Ground Truth Results
In order to study the isolated performance of multi-task models, we provide part ground truth information in advance during the inference stage. Specifically, we provide ground truth target texts for post-explaining and report their performance of explanation. Conversely, we offer ground truth explanations for pre-explaining and report their performance of correction. This experimental setting allows for revealing the specialized performance, eliminating the distraction of previously generated contents. The results are presented in Table 4.
For the task of explanation, two base models slightly outperform chat models.
Specially, the base version models of Qwen and DeepSeek exhibit a minor increase in performance over their chat/instruct counterparts on classifying error types and providing error descriptions. However, this is not true for Llama3, where the Llama3-instruct model obtain the highest Acc, METEOR and ROUGE scores. Also noteworthy is the significantly lower miss rates of chat/instruct models compared to base models, indicating a tendency for the latter to overlook explanations, even when ground truth target texts are available. These findings contradict the joint results in Table 3. We speculate the reason is base models may be more susceptible to low-quality self-generated corrections.
Ground truth explanations tremendously improve correction performance.
Since the explanations include explicit clues for corrections such as evidence words and revision advice, it is effortless for pre-explaining models to correct the source.
6.4 Comparison with Pipeline
We compare the results of multi-task models and GEC-GEE pipeline with COTE in Table 5. It indicates that the pipeline can improve both correction and explanation performance compared to multi-task models, highlighting the challenges of learning a multi-task model for EXGEC. A significant advantage of the pipeline solution is its superior performance of the correction task, which can subsequently enhance the explanation performance. However, adopting the pipeline solution requires heavy deployment and extra training costs.
7 Analysis
7.1 Ablation Results
We conduct ablation studies on Qwen1.5-7B-chat to provide in-depth insights into post-explaining models. We also study the effect of model sizes in Appendix B.2 and provide a case study for different LLMs in Appendix B.3.
[…]
w COTE 99.93 0.43 81.53 39.56 0.74 17.88 41.40 51.73 / 25.81 / 36.51
w/o COTE 49.64 54.01 42.51 17.77 0.93 11.53 33.81 46.35 / 19.34 / 31.28
Table 7: Ablation results of COTE from the same Qwen1.5-7B-chat post-explaining model.
@@ 7.1 Ablation Results
Table 8: Correlations between human judgements and metrics for error descriptions.
Effect of COTE.
The impact of COTE introduced in Section 5.2 is examined in this section. We provide the post-explaining model with ground truth target texts, which allows us to focus on the explanation performance. The results presented in Table 7 reveal a huge performance drop if we do not leverage COTE, especially the hit and miss rates. This demonstrates the effectiveness of COTE.
Effect of token-wise decoding strategies.
By default, we employ beam search decoding for corrections and top-p decoding for explanations. In this section, we explore the reverse setting, and the results are reported in Table 6. When switching from beam search to top-p for correction, we observe a huge performance drop in precision and F0.5 and increase in recall, which means top-p encourages LLMs to over-correct Cao et al. (2023). On the other hand, leveraging beam search improves explanation performance, suggesting the potential benefits of a greedy decoding algorithm for the task. However, we notice that beam search also increases the miss rate. We speculate that beam search may discard some low-likelihood explanations.
7.2 Human Evaluation for Error Descriptions
@@
Despite the efficiency of automatic metrics in evaluating error descriptions, their accuracy remains to be confirmed. Therefore, this section attempts to demonstrate the suitability of different metrics by comparing their corrections with human judgements. We report the correlations between two human annotators and the ones between average human ratings and metric scores in Table 8. We observe the inter-annotator correlations are close to 1, meaning it is relatively easy to determine the quality of error descriptions for human. Among various metrics, ROUGE-1 achieve the highest correlations, followed by METEOR. All the introduced metrics show moderate or high correlations, indicating that it is advisable to employ them as proxies for human evaluation. We provide the detailed annotation guidance and rating rules in Appendix B.4.
8 Conclusion
@@
We propose and formulate the task of EXGEC, overcoming the limitation of previous studies that fail to establish the interaction of both correction and explanation tasks. To develop the task, we propose the EXCGEC benchmark, based on which we develop baseline models in multiple settings. Extensive experiments and analyses reveal several challenges of the task, and we hope this paper can serve as a starting point for future exploration.
Limitations
Inferior performance of multi-task models.
In our experiments, we observe the pipeline solution outperform the multi-task solutions, regardless of correction or explanation tasks. This suggests that the multi-task models struggle to reap positive benefits from the interaction of both tasks. We leave the exploration of effective multi-task learning EXGEC models to the future work.
Limitations of synthesizing datasets.
LLM-augmented datasets may include some unintended biases towards or inaccuracies, resulting in skewed or unfair outcomes in applications. Second, it is necessary to manually filter out invalid data in order to ensure the quality of datasets. But it is indeed a advisable method to construct datasets using LLMs, considering its efficiency.
Adaptation to other languages.
The general design of our proposed edit-wise explanations can be easily adapted to other languages. However, the detailed design may not be suitable to other languages. For example, the two-tier hierarchical taxonomy of error types is tailored for Chinese.
Ethics Statement
Our proposed benchmark is built upon existing datasets, backbones and metrics, all of which are publicly available. We have cited the corresponding authors or projects of them, and confirm that they are consistent with their intended use.
Additionally, we conduct human evaluation experiments to ensure the quality of the dataset and find out the correlations between metrics and human judgements. To achieve this, we hire 12 native speakers, all of whom are graduated students. Each annotator could complete the entire annotation process within approximately 6$\sim$8 working hours. All annotators were paid for their work, with an average salary of approximately \$5 per hour.
Appendix A Benchmark Details
A.1 Description of Grammatical Error Types
In the taxonomy of Chinese grammatical errors, we first divide grammatical errors by their effecting granular into 5 major types, including punctuation-level, spelling-level, word-level, sentence-level, and other special errors. This section will clarify all involved grammatical error types by providing detailed descriptions and corresponding examples listed in Figure 456.
Punctuation-level Error.
This type of grammatical errors primarily involves redundancy, missing, and misuse of punctuation.
  • 标点冗余 (Punctuation Redundancy). The type of punctuation redundancy refers to the insertion of punctuation in unnecessary places. For punctuation redundancy errors, first explain the role of the punctuation symbols involved, and then explain the reasons for punctuation redundancy in the current case.
  • 标点丢失 (Punctuation Missing). Punctuation missing mainly refers to the omission of punctuation that should have existed in the middle and end of a sentence. For the explanation of punctuation missing errors, first point out the evidence words and missing punctuation symbols, and then explain the role of the added punctuation in this context.
  • 标点误用 (Punctuation Misuse). Misuse of punctuation is very common in daily Chinese writing. For cases of punctuation misuse, first briefly explain the roles of misused punctuation and correct punctuation, and then explain sufficient reasons for correction.
Spelling-level Error.
Spelling-level errors refer to people who, due to carelessness or lack of knowledge, write incorrect characters or words during the writing process. The type is so common that Chinese Spelling Check (CSC), as a standard NLP task specialized in spelling-level errors, attracts the attention of many researchers. Inspired by these studies, we categorize spelling-level errors further into 4 sub-classes.
  • 字音混淆错误 (Phonetic Confusion Error). Phonetic confusion errors are caused by misusing Chinese characters with the same or similar pinyin. The vast majority of Chinese Internet users use the pinyin input method, so many Chinese spelling-level errors on the Internet fall into this type.
  • 字形混淆错误 (Glyph Confusion Error). In addition to the pinyin input method, some users apply the Wubi input method or other glyph-based input methods. In this case, they are prone to spelling errors due to confusion of fonts or strokes.
  • 词内部字符异位错误 (Internal Character Misplacement Error). Internal character misplacement error refers to expressing a multi-character word in the disorder of characters. The type seldom happens for native speakers, but sometimes in texts written by L2 speakers. For example, the spelling-level error “共公” falls in this type and should be corrected to “公共”.
  • 命名实体拼写错误 (Named Entity Misspelling). There are numerous named entity words in Chinese, such as person names, organization names, place names, and all other entities identified by terminologies. These words are also very prone to spelling errors.
Word-level Error.
Word-level errors often refer to misuse of individual words or idioms in a sentence, but the syntactic structure of the sentence is correct. This type of error belongs to the most common category in Chinese text errors and can usually be subdivided into the following three types:
  • 词语冗余 (Word Redundancy). The simultaneous appearance of words with the same or similar meanings in a sentence can cause semantic repetition and sentence redundancy, which is known as word redundancy. Repeated words often appear adjacent to each other, so it is important to pay attention to whether the meanings of adjacent words are exactly the same. If they are the same, it may lead to the problem of word redundancy.
  • 词语丢失 (Word Missing). In modern Chinese, sentences generally have six major components, namely subject, predicate, object, attributive, adverbial, complement, etc. A sentence must express a complete meaning, and its structure must also be complete. The so-called complete structure does not mean that a sentence must have the usual six components, but rather that the sentence should be composed of the necessary components to express the complete meaning. If the necessary sentence components are missing, it will cause the phenomenon of word missing.
  • 词语误用 (Word Misuse). Word Misuse indicates improper use of words in the text. The main cause of this error is the author’s insufficient understanding of the meaning and part of speech of a certain word.
Sentence-level Error.
This type mainly involves sentence-level issues, not just individual words or characters. Sentence-level errors are often caused by violating common syntactic structures, or not following objective reasoning.
  • 词序不当 (Improper Word Order). Proper word order is essential to express exact meaning in Chinese. Writing texts without accurate word order results in the type of improper word order. If a sentence is not combined according to the intended meaning, it may lead to confusion in the sentence structure, resulting in an imbalance in the relationship between sentence components and affecting the expression of sentence meaning.
  • 逻辑不通 (Illogicality). Illogicality refers to a sentence that conforms to grammatical norms but does not conform to logical reasoning. Illogicality can be caused by many reasons such as improper logical order, causal confusion, and reversal of subject and object.
  • 句式杂糅 (Run-on Sentence). Run-on Sentence in Chinese usually refers to the use of two formats or sentences with similar or identical meanings in one sentence. People originally used one format when writing sentences, but due to interference from other factors such as sentence content, they may unconsciously switch to another format, resulting in a mixture of the two formats.
Other Special Error.
Besides the above grammatical error types, other several types can not easily fit in the mentioned major types. So we classify them to other special errors.
  • 照应错误 (Inconsistency Error). Inconsistency errors are ones involved in the mistaken referential relationship between two words, and explaining this grammatical error requires knowledge of the referential relationship between each word.
  • 歧义错误 (Ambiguity Error). Ambiguity errors happen when a word or a sentence can be understood as having multiple meanings.
  • 语气不协调 (Inconsistent Tone). Inconsistent tone refers to the inconsistency of tone between the preceding and following sentences.
Additionally, we define the grammatical error type Other as ones that do not fit in any of the above error types. These errors are usually involved in rather significant modification and sometimes change the original semantics.
Figure 3: Distribution of 7 kinds of LLM errors.
Configuration Value
Fine-tuning
Devices 2 Tesla A100 GPU (80GB)
Epochs 5
Finetuning type Lora
Train batch size per GPU 2
Eval batch size per GPU 1
Gradient accumulation steps 16
Optimizer
AdamW
($\beta_{1}=0.9,\beta_{2}=0.98,\epsilon=1\times 10^{-6}$)
Learning rate $5\times 10^{-5}$
Learning rate schedule cosine decay
Warmup steps 20
Eval steps 200
Cutoff length 1024
Preprocessing workers number 16
Numerical precision fp16
Weight decay 0.05
Inference
Beam size 5
Top-p 0.8
Max new tokens 2048
Temperature 0.7
Table 9: Hyper-parameters used in our experiments.
[…]
Qwen1.5-1.8B-chat 21.11 / 19.28 / 20.72 28.91 / 15.70 / 24.74 59.94 65.14 55.80 23.27 0.89 10.19 34.35 48.66 / 22.70 / 34.23
Qwen1.5-4B-chat 22.49 / 20.84 / 22.14 30.57 / 16.85 / 26.29 62.91 62.70 57.16 25.31 0.85 11.61 35.91 46.83 / 19.59 / 30.86
Table 10: Comparison of post-explaining models with various model sizes.
Figure 4: Examples of error types.
Figure 5: Examples of error types.
Figure 6: Examples of error types.
A.2 Examples of Error Types
We list the examples of error types in Figure 456.
A.3 Prompt of Generating Explanations
The prompt we use to generate explanations is shown in Figure 8. We also provide an English version in Figure 9.
A.4 Detailed Description of LLM Failure Modes
We categorize the failure modes in our case into seven major reasons: incorrect type, incorrect severity, incorrect format, incorrect template, non-fluency, unreasonability, and unfaithfulness. One expert annotator is asked to classified the sampled 100 invalid explanations, where an explanation may be categorized into multiple failure modes. The annotation results, illustrated in Figure 3, reveal that GPT-4 tend to mis-classify grammatical errors and providing unfailthful error descriptions. On the other hand, GPT-4 is capable to a large extend to offer well formed, fluent, and reasonable explanations, demonstrating the effectiveness of LLM annotation on this task.
The definitions of seven failure modes of explanations are as follows:
  • Incorrect type: the error type is incorrect.
  • Incorrect format: the evidence content and the correction content are not highlighted by special markers 【 】 or { }.
  • Incorrect template: the error description does not follow the syllogism form of deductive reasoning.
  • Non-fluency: the error description is non-fluent or unreadable.
  • Unreasonability: the error description contains obvious mistakes about linguistics, thus making it unacceptable for human.
  • Unfaithfulness: the error description is not targeted to the given edit.
Appendix B Experimental Details and Extra Results
B.1 Implementation Details.
We train all models for 5 epochs and select the best model validated on EXCGEC-dev and report its performance on EXCGEC-test. The detailed training hyperparameter values of the all models in our experiments are shown in Table 9.
B.2 Effect of Model Sizes
Table 10 indicates the varying performance across model sizes ranging from 1.8B to 7B. We observe consistent performance enhancement with increasing model sizes.
B.3 Case Study
We provide a case study in Table 7.
B.4 Details of Human Rating
Specifically, we hire 2 native Chinese speakers to rate the explanations generated by 6 post-explaining models in Table 3 conditioned on ground truth target texts. The rating scores range from 0 to 100, and each annotator concurrently rate 6 explanations for each sample. We randomly select 100 samples for annotation. We provide annotators with general scoring suggestions:
  • 100 points: Explain and describe fluently (fluency), introduce relevant semantic knowledge to enhance persuasiveness (rationality), and explain that it is aimed at the current editor (loyalty). All aspects are impeccable, and there is almost no better explanation or description than this.
  • 80$\sim$100 points: The explanation and description are expressed fluently, satisfy fidelity, and have a certain degree of rationality, but there are certain degrees of flaws.
  • 60$\sim$80 points: The explanation and description are expressed fluently, but the fidelity or rationality is not good enough, but it is somewhat helpful for correcting the grammar error in understanding.
  • 30$\sim$60 points: The explanation and description are expressed fluently, but the rationality is poor, and it is not very helpful for correcting the grammar error in understanding.
  • 0$\sim$30 points: The explanation and description are expressed fluently, but the fidelity is poor, and the object of explanation is not the current editor. There is no help in correcting the grammar error for understanding.
  • 0$\sim$30 points: The explanation and description are vague and cannot be understood. There is no help in correcting the grammar error for understanding.
Figure 7: A case study of all the LLMs involved in our experiments.
Figure 8: The prompt used for explanation generation. For each error type, We provide the definition, a suggested template of error description, and a demonstration for GPT-4.
Figure 9: The English prompt used for explanation generation.

### v2

@@ Introduction
Inference.
For post-explaining EXGEC models, we design a specific Correct-Then-Explain decoding algorithm called COTE, which is presented in Algorithm 1. First, we employ the greedy beam search decoding strategy for the correction part, which is beneficial to relieve the over-correction problem that is common in LLMs. Then, we apply CLEME to extract edits. Notably, we merge adjacent edits with a distance of less than 2 characters to avoid fragmented edits. Finally, we leverage the Top-p decoding strategy for generating explanations, encouraging diversified natural language explanations. It is worth noting that COTE is not accessible to pre-explaining models since the edit extraction tool necessitates both a source text and a target text.
Experiments
@@
We benchmark mainstream LLMs including Qwen-1.5 (Bai et al. 2023), Llama-3 (Touvron et al. 2023), and DeepSeek (Bi et al. 2024). For these LLMs, we experiment with their base and chat (or instruct) versions to investigate whether further alignment training benefits the task. All experimental results are averaged over three runs with different random seeds on EXCGEC-test in Table 2. More training details are reported in the supplementary materials.
@@
We obtain the metric results using public toolkits including ROUGE (Lin 2004), NLTK (Bird and Loper 2004), and scikit-learn (Pedregosa et al. 2011). Particularly, we observe many hypothesis edits are not covered by the corresponding reference edits, making it impossible to subsequently evaluate the explanations for these edits. To address this, we introduce two extra indicators, namely Hit and Miss rates. A hypothesis edit overlapping with a reference edit is designated as a hit edit, while a reference edit without any match with hypothesis edits is deemed a miss edit. The hit rate is defined as the ratio of hit edits to all hypothesis edits, and the miss rate as the ratio of miss edits to all reference edits. Only the hit edits are used to calculate the evaluation outcomes for explanations.
Results of Multi-task Models
Table 3 presents the main results of multi-task models.
Post-explaining models outperform pre-explaining models.
Concerning the correction aspect, all post-explaining models consistently obtain higher F0.5 scores than pre-explaining models, regardless of the applied backbones. A similar pattern is observed in the explanation part, where all the pre-explaining models invariably underperform their post-explaining counterparts. This suggests complexity for LLMs to directly explain grammatical errors without auxiliary information like target sentences or extracted edits. And once pre-explaining models generate flawed explanations, the ensuing distraction impedes their ability to accurately correct the source text.
Chat models are superior to base models.
For post-explaining models, we observe all chat or instruct models gain slightly higher F0.5 correction scores, and they also marginally outperform their base version counterparts in the explanation task. It indicates that additional alignment training (Wang et al. 2023) can benefit the EXGEC task.
@@
Ground Truth Results
To examine the isolated performance of multi-task models, we introduce partial ground truth information in advance during the formal inference stage. This is achieved by pre-inserting ground truth corrections or explanations into the decoding phase prior to formal inference. Specifically, we utilize ground truth target texts for post-explaining and evaluate the performance of the explanation task. Conversely, we provide ground truth explanations for pre-explaining and assess the performance of the correction task. This approach enables a detailed analysis of each task’s performance under oracle conditions. The results, as depicted in Table 4, reveal that the incorporation of ground truth information significantly enhances performance. Notably, post-explanatory models equipped with ground truth corrections exhibit a marked improvement in explanatory performance across all LLMs. This observation extends to post-explanatory models with ground truth explanations, suggesting that previously generated low-quality content adversely affects subsequent generative processes.
Comparison with Pipeline
We compare multi-task models and a GEC-GEE pipeline with COTE in Table 5. It indicates that the pipeline solution can improve both the correction and the explanation performance compared to multi-task models, highlighting the challenges of learning multi-task models for EXCGEC. However, adopting the pipeline solution requires heavy deployment and training costs. We speculate that LLMs with only 7B parameters cannot establish intimate interaction of correction and explanation tasks.
Analysis
Ablation Results
We conduct ablation studies on Qwen1.5-7B-chat to provide in-depth insights into post-explaining models. We also study the effect of model sizes and provide a case study for different LLMs in the supplementary materials.
@@
w COTE 99.93 0.43 81.53 39.56 0.74 51.73 / 25.81 / 36.51
w/o COTE 49.64 54.01 42.51 17.77 0.93 46.35 / 19.34 / 31.28
Table 6: Ablation results of COTE from the same model.
Effect of COTE.
We introduce COTE that provides gold alignment for post-explaining models, thus unburdening LLMs during the inference stage. The impact of COTE is quantitatively examined in this section. We provide the post-explaining model with ground truth target texts, which allows us to focus on the explanation performance. The results presented in Table 6 reveal a huge performance drop if we do not leverage COTE, especially the hit rate and the miss rate. This demonstrates the effectiveness of COTE.
Human Evaluation for Error Descriptions
@@
Table 7: Correlations between human judgements (A1 and A2) and metrics results for error descriptions.
We adopt traditional metrics for assessing the quality of generated error descriptions mainly for their reproductivity and efficiency (Clinciu, Eshghi, and Hastie 2021). However, their reliability requires further validation. Therefore, this section attempts to demonstrate the suitability of these metrics through their corrections with human judgments. We assign two human annotators to score the error descriptions generated by all 6 post-explaining models, with the scoring scale from 0 to 100. For each sample, the annotators are instructed to concurrently evaluate all the error descriptions, referencing a gold explanation generated by GPT-4 to guarantee a rigorous and reliable assessment. Additional details are delineated in the supplementary materials.
@@
We report Pearson and Spearson correlations between the metric results and the human judgments in Table 7. We observe the inter-annotator correlations are close to 1, meaning it is relatively easy to determine the quality of error descriptions for human annotators. Most metrics achieve moderate or high correlations with human judgments, which means that it is relatively reasonable to use simple n-grams-based metrics to evaluate the quality of error descriptions efficiently. Among various metrics, ROUGE-1 achieves the highest correlations, followed by METEOR. All the introduced metrics show moderate or high correlations, indicating that it is advisable to employ them as proxies for human evaluation. We provide detailed annotation guidance and rating rules in the supplementary materials.
Conclusion
@@
We propose and formulate the task of EXGEC, establishing the interaction of correction and explanation tasks. To develop the task, we propose the EXCGEC benchmark, based on which we build baseline models. Extensive experiments and analyses reveal several challenges of the task. We hope this paper can serve as a starting point for future exploration.
Acknowledgements
This research is supported by National Natural Science Foundation of China (Grant No.62276154), Research Center for Computer Network (Shenzhen) Ministry of Education, the Natural Science Foundation of Guangdong Province (Grant No.2023A1515012914 and 440300241033100801770), Basic Research Fund of Shenzhen City (Grant No.JCYJ20210324120012033 and GJHZ20240218113603006), the Major Key Project of PCL for Experiments and Applications (PCL2021A06).